### v5

# Chemist-X: Large Language Model-empowered Agent for Reaction Condition Recommendation in Chemical Synthesis

Abstract
Recent AI research plots a promising future of automatic chemical reactions within the chemistry society. This study proposes Chemist-X, a transformative AI agent that automates the reaction condition recommendation (RCR) task in chemical synthesis with retrieval-augmented generation (RAG) technology. To emulate expert chemists’ strategies when solving RCR tasks, Chemist-X utilizes advanced RAG schemes to interrogate online molecular databases and distill critical data from the latest literature database. Further, the agent leverages state-of-the-art computer-aided design (CAD) tools with a large language model (LLM) supervised programming interface. With the ability to utilize updated chemical knowledge and CAD tools, our agent significantly outperforms conventional synthesis AIs confined to the fixed knowledge within its training data. Chemist-X considerably reduces chemists’ workload and allows them to focus on more fundamental and creative problems, thereby bringing closer computational techniques and chemical research and making a remarkable leap toward harnessing AI’s full capabilities in scientific discovery.
keywords:
Retrieval-augmented generation, computer-aided synthesis, large language models, reaction condition recommendation, AI for chemistry
1 Introduction
Recent “artificial intelligence (AI) for chemistry” research aims to free human labor with AI-supervised chemical reaction platforms [1, 2, 3, 4]. These innovative platforms employ machine learning (ML) algorithms and data-driven methods to orchestrate synthetic routes and identify optimal reaction conditions for maximizing yields. In the foreseeable future, chemical robots equipped with integrated AIs will autonomously design and refine chemical reactions, thus allowing human experts to focus on more creative and foundational research. Yet, traditional AI systems fall short of human experts who possess a broad spectrum of chemical knowledge and can continually update their understanding by consulting related literature when dealing with new problems. Conventional AI models, being confined to the knowledge from their training data, exhibit limited generalization ability when encountering unknown reactions. They can provide chemists with preliminary assistance for familiar reactions but cannot realize the ambitious aim of fully automated synthesis of unprecedented products.
Recent technical breakthroughs in retrieval-augmented generative AI (RAG-AI) have shed light on this problem. RAG-AI, as its name suggests, harnesses the strengths of retrieval-based methods and the capacities of generative AIs like large language models (LLMs). RAG enables an AI to retrieve relevant information from a vast and continuously updated knowledge source [5, 6], while LLM equips the AI with close-to-human abilities in information analysis and code writing [7, 8]. This paper develops a sophisticated chemical reaction agent, namely Chemist-X, with RAG-AI technology. As a proof of concept, we focus on RAG for reaction condition recommendation (RCR), which is equally important as retrosynthesis but receives less attention in chemical synthetic research.
LLMs like ChatGPT [9] or GPT-4 [10], thanks to their advanced generalization capabilities and near-human intelligence, demonstrate remarkable proficiency in devising feasible execution plans for RCR problems. The plan proposed by the LLM can be summarized as the following methodical three-phase approach: 1) identify structurally analogous molecules, which may share similar chemical characteristics with the target products, to provide foundational references for further analysis; 2) review existing literature about these analogous molecules to narrow down the chemical space of potential reaction conditions; and 3) utilize Computer-Aided Design (CAD) technology to select a subset of promising conditions from the refined chemical space. This systematic ”search-analyze-recommend” pattern suggested by the LLM well aligns with the methodology a human expert might undertake when addressing the RCR challenge associated with an unfamiliar molecule, thereby providing strong evidence for the practicality of the LLM’s three-phase solution. Fig. 1 illustrates the three-phase solution with more technical details.
Figure 1: The three-phase RCR framework of Chemist-X. a) The agent’s problem-solving procedure presented in the form of a pipeline. All three phases are automatically executed with the control Chemist-X. b) A detailed illustration of operations involved in Chemist-X. The LLM is the kernel of the agent. It is important to note that the figure follows the recent concept of “AI agent”, in which an LLM-empowered AI system can achieve close-to-human performance in a specific task owing to its abilities in i) using additional knowledge as “memory”, ii) leveraging “tools” like pre-defined APIs and auto-generated codes, and iii) taking necessary actions so that the processing can automatically move on to the next step.
Phase One retrieves molecules analogous to the target product by leveraging an LLM-empowered code generation scheme tailored to application programming interfaces (APIs) provided by molecular databases such as PubChem [11] and ChemSpider [12]. These databases typically offer APIs to facilitate information retrieval and provide examples of usage in their tutorial handbooks. However, software programming based on API documents requires a related background in computer science, and it could be a time-consuming task given that each database has its own API protocols and documentation. Phase One makes such API-based information retrieval accessible to researchers regardless of their familiarity with software programming, thereby bridging the gap between chemists and computer scientists.
In pursuit of automatic API programming without human intervention, we harness in-context learning (ICL) [13, 14], an effective few-shot learning method that equips the LLM with necessary API knowledge as background information and augments its code generation capabilities using exemplar codes in prompts. It is important to note that providing the LLM with the entire programming handbook in ICL is counterproductive, as this could overwhelm the LLM, making it struggle to pinpoint the most relevant code segments. To address this problem, we put forth the concept of top match slice (TMS), which refers to the portion of the document bearing the utmost semantic similarity to the problem description provided to the LLM. This similarity is measured by the distance between our problem description and the tested document slice within the semantic embedding space [15]. Our experiments indicate that incorporating the TMS of an API handbook into the ICL prompt significantly enhances the LLM’s performance in retrieving information about similar molecules via APIs. The scheme outperforms ICL with complete documentation or random document slices in multiple aspects, including higher code reliability, better cost-efficiency, and reduced response time.
Phase Two retrieves detailed literature information from online sources and analyzes the data with automatically generated code. Upon the completion of Phase One, we possess a list of structurally similar molecules. To understand how previous papers deal with similar reactions, our agent first searches for chemical reactions related to each of these analogous molecules in online literature platforms like PubMed [16] via web crawlers. With HTML content collected by these crawlers, the agent then distills the raw data to construct a potential reaction condition subspace for the target molecules.
Beyond the creation of the internet crawler, Chemist-X also realized an efficient HTML analyzing scheme that overcomes the limitation of the LLM’s input information length. An online literature platform may employ complex web designs, making its HTML source surpass the maximum token length permissible for LLMs. Chemist-X addresses this obstacle by instructing the LLM to generate Python code tailored to HTML analysis, rather than directly feeding the raw HTML source to the LLM as in prior works [17, 18]. Compared with the conventional scheme, our script-based method demonstrates higher accuracy and efficiency in extracting reaction condition data from extensive HTML sources.
Phase Three makes final recommendations. Once the chemical reaction subspace has been refined, our focus shifts to identifying the optimal set of reaction conditions with top yield expectations. There are some prior studies investigating CAD schemes for such recommendation tasks [19, 20, 21]. Given the intense research interest in RCR for chemical synthesis, there may be more advanced CAD schemes in the future. Chemist-X serves as a universal platform for cutting-edge CAD algorithms, wherein the latest research findings could be integrated in the form of APIs. To fulfill this objective, our agent harnesses LLMs to comprehend the functionality of each chemical CAD algorithm via its documentation and to select appropriate CAD tools through API programming. As the chemistry community continues to update CAD tools, our platform enables the continuous evolution of Phase Three, ensuring it remains at the forefront of technological advancement.
This paper developed an advanced ML model for RCR tasks and packaged it into an API so that the agent can execute the new model for concept proving. Specifically, in our new model, we designed a novel chemical reaction fingerprint and applied it to enhance the conventional ML models. The fingerprint diverges from traditional reaction encoding approaches in two key aspects. First, we apply chemistry-informed molecular graph (CIMG) [22], an innovative molecular descriptor based on graph neural networks (GNNs) [23], as the underlying molecule encoding method to boost the quality of molecular descriptions. Second, we integrate supervised contrastive learning (SCL) [24] in our fingerprint-generation network, thus enhancing the model’s ability to distinguish the unique features of high-yield reactions from others. Further, to better serve the SCL process, we replaced the noise-sensitivity numerical descriptions of reaction yields with coarse categorical yield labels. With the above efforts, our reaction embedding scheme, denoted as the coarse-label SCL (CL-SCL) fingerprint, demonstrates exceptional efficacy in capturing the essential characteristics of high-yield reactions. It constantly outperforms prior studies in terms of precision and robustness across a variety of ML models.
In general, the three-phase framework supervised by LLM, along with the specific techniques employed in each phase, enables Chemist-X to automatically acquire the latest online molecules/literature knowledge and to utilize the most cutting-edge CAD tools. With the advantages of continuous evolution and unmanned processing, Chemist-X makes a significant contribution to the AI-chemistry community and makes a meaningful step towards the ultimate goal of AI-supervised chemical reaction platforms.
2 Results and Discussions
This section demonstrates how Chemist-X completes the RCR task with an example Suzuki reaction targeting on 6-(1-methyl-1H-indazol-4-yl) quinoline, whose SMILES is written as CN1C2=CC=CC(C3=CC=C4N=CC=CC4=C3)=C2C=N1. We investigate the agent’s performance in each phase via comprehensive unit tests, and then we conduct wet-lab experiments to validate the recommendation results.
2.1 Unit Tests
2.1.1 Unit Tests for Phase One
Phase One employs ICL to craft Python code capable of interfacing with PubChem via its official APIs. Figure 2a illustrates our prompt, which includes a fixed part and a flexible part. Our question is described in the fixed part, while the reference example of ICL should be contained in the flexible part. Our agent identifies the most pertinent section of the API documentation and uses it to fill in the flexible part of the prompt.
During our experiment, Chemist-X splits the text of PubChem’s API documentation [11] into eight discrete slices with similar lengths. It then employs text-embedding-ada-002 (henceforth referred to as ADA-002), an advanced natural language processing (NLP) model from OpenAI [25], to convert these document slices and our query prompt into vector representations within a semantic space. After semantic embedding, we measure the semantic similarity between the vector of each document slice and that of our prompt. For a comprehensive evaluation, we test two classical semantic similarity metrics in NLP research: Cosine Similarity [26] and L2 Similarity [27]. For a detailed explanation of the embedding procedure and the two metrics tested, we refer readers to the methodology section.
If a slice exhibits high semantic similarity to our questioning prompt, it is likely relevant to our task and could be considered a potential candidate for completing the missing part of the prompt shown in Figure 2a. Our experimental findings, depicted in 2b and 2c, confirm this argument: across all eight candidates, slice #3 achieves the highest semantic similarity regardless of the metric applied. Upon manually reviewing PubChem’s documentation, this slice indeed pertains to the “similar molecules inquiry function” provided by PubChem and includes a practical example that guides users on how to implement the inquiry function through programming. Therefore, the agent treats slice #3 as the TMS and uses it to fill in the prompt, eventually making the LLM perform more effective ICL.
The agent’s accuracy and efficiency are substantially improved by incorporating the TMS. We compare our TMS-ICL approach against three alternatives: (i) zero-shot learning without any background information or examples, (ii) ICL using the complete API documentation as background reference, and (iii) ICL with a randomly chosen slice as background reference. We conduct ten trials of our TMS-ICL and the three alternative methods using a GPT-4 model. We assess the performance of these four code-generation strategies based on computational resources used and the success rate in generating functionally correct code in the ten trials. Specifically, we consider four aspects of computational requirements: the average number of question/answer tokens used in one interaction with the GPT-4 model, the model’s average response time in completing one code design task, and the average cost of utilizing the commercial LLM model for one time. Experimental results presented from Figure 2d to Figure 2h indicate that our TMS-ICL method consumes slightly more computational resources than the random slice approach yet delivers a remarkably higher success rate than all three alternatives. The zero-shot approach uses the least computation resources but has a disappointingly low success rate. The all-document approach shows modest improvement over zero-shot learning but does not match the performance of our TMS-ICL and requires substantially more computational resources.
Figure 2: Experimental details of Phase One’s unit test. a) The prompt we used for the auto code generation in this experiment. Here <>expectation<...>< … > is a placeholder representing the flexible part of the prompt to be selected with the TMS scheme. b) and c) Similarity scores between each documentation slice and the fixed part of our prompt. The scores in b) and c) are calculated with cosine distance and L2 distance, respectively. d) to h) Compression of four different prompting schemes (zero-shot/all-document/random-slice/TMS prompting) in terms of cost and performance.
2.1.2 Unit Tests for Phase Two
Phase Two retrieves reaction-related information from literature platforms and extracts chemical reaction condition data, such as bases and solvents, from HTML sources. We develop an innovative prompting framework for generating web crawlers via ICL (see 3a) and an HTML analysis scheme that employs LLMs to generate Python code tailored to information extraction from HTML contents (see 3b).
We apply GPT-4 in generating the web crawler. We note that GPT-4 offers customized parameter configuration for users’ flexible adaption. Among these parameters, temperature and presence penalty play a vital role in code generation: the former affects the creativeness of the output, while the latter influences the repetition of generated content. In Figure 3a, we investigate how these two parameters affect the model’s performance in our case. As we can see, our task is sensitive to the model’s temperature and performs well when the temperature is low, given that being focused and deterministic is more important than creative thinking in our web crawling task.
As for HTML analysis, experiments presented in Figure 3d demonstrate that our approach significantly outperforms conventional HTML analysis methods that input the entire HTML source into LLMs directly [17, 18]. Our experiments focus on the top 30 structurally similar molecules identified in Phase One. The HTML source of each molecule’s search result has been collected by the crawler. Upon manual review, these files include 143 pieces of reaction condition data in total. The conventional approach requires 30 separate interactions with the GPT-4 model, i.e., one round of interaction for each HTML file; while our method involves a single interaction for code generation, which is then repeatedly executed to extract information from these 30 files. Experimental results prove that our scheme consumes significantly fewer question/answer tokens than the conventional scheme in the 30-HTML task, which further results in fewer costs and shorter response time. Despite consuming fewer computational resources, our method achieves superior extraction accuracy, successfully identifying in total 141 out of 143 pieces of correct data from these HTML sources, while the conventional scheme only figures out 78 pieces of data. Finally, we calculate the F1 score of the two schemes: 70.59% for the conventional scheme and 99.30% for our scheme.
Figure 3: Comparative Analysis of HTML Extraction Methods. a) The prompt we used for developing the web crawler. b) The prompt we used for developing the HTML analyzer. Here we apply the ICL method with three examples and hint the LLM to apply a character-matching strategy when generating the code. c) A 3D figure showing how the generation process of the web crawler is affected by the LLM’s parameter. d) A benchmarking between our code-based HTML analysis method with the conventional method that inputs HTML directly into the LLM. Notably, to make the comparison possible, we simplified the challenge of the conventional scheme by pre-processing the raw HTML to remove extraneous content, such as formatting and search bar. Otherwise, the length of some raw HTML sources would exceed GPT-4’s token limit. Despite this adjustment, our method still demonstrates higher precision and lower resource usage.
2.1.3 Unit Tests for Phase Three
Phase Three develops a novel RCR algorithm with CL-SCL fingerprint as a proof of concept to show Chemist-X’s ability to utilize updated CAD tools. Unit tests in this subsection focus on the efficacy of the proposed CAD scheme, while a systematic validation of the agent is presented in the next subsection via web-lab experiments. To examine the generalization ability and the universality of the fingerprint, experiments are conducted as follows: we use a Buchwald-Hartwig reaction dataset [20] for model training and test the model with a Suzuki-Miyaura reaction dataset [28]. Further, in the evaluation, we follow the traditions established in prior RCR works [21] and use the maximum observed yield as the figure of merit. We denote the maximum observed yield of final recommendations by $\mu_{N}$, where $N$ is the batch size (BS) of the final recommendation.
Benchmarking Experiments: We first compare our CL-SCL fingerprint against two reaction fingerprints from prior studies: DRFP [29] and Mordred [30]. We validate these fingerprints with three ML algorithms, i.e., random forest (RF) [31], tabular transformer (TT) [9], and Xgboost (XGB) [32], under different BS configurations. Table 1a presents the resulting $\mu_{N}$ across different experimental setups, which reveal that our CL-SCL fingerprint consistently delivers superior reaction condition recommendations than other alternatives regardless of BS setting and the ML model harnessed.
DRFP Fingerprint Mordred Fingerprint Our CL-SCL Fingerprint
BS TT RF XGB TT RF XGB TT RF XGB
1 39.242 53.950 52.750 48.876 34.016 48.118 61.253 56.510 53.553
2 40.290 63.676 68.907 53.716 40.711 63.778 75.673 68.087 68.937
3 70.258 72.153 72.970 65.806 77.386 77.796 86.953 79.663 70.653
4 86.365 80.251 76.790 75.998 81.474 83.390 90.437 83.109 74.223
5 89.346 88.280 77.400 85.534 83.588 86.327 91.363 86.551 79.665
(a) Experimental Results of Benchmarking Experiments ($\mu_{N}$ is recorded in the form of percentage)
i. Remove the CL-SCL network ii. Remove CIMG encoding iii. Our original method
TT 80.232 82.273 91.363
RF 81.186 76.140 86.551
XGB 73.899 76.955 79.665
(b) Experimental Results of Ablation Study ($\mu_{5}$ is recorded in the form of percentage)
Table 1: A detailed result record for experiments conducted in Phase Three.
Ablation Study: There are two critical operations in generating the fingerprint: 1) the CIMG molecule encoder and 2) the CL-SCL network. A question one may ask is the effectiveness and necessity of each operation. To address this query, we conducted an ablation study that tests the indispensability of each element. The ablation study sets the BS as 5. In Experiment (i), we remove the CL-SCL network, allowing the chemical reactions encoded by CIMG to directly interface with subsequent ML models. In Experiment (ii), we substitute the CIMG encoder with a trivial encoder that simply concatenates molecule encodings. Beyond these alterations, the rest experimental setting mirrors the benchmarking experiment.
Table 1b presents the resulting $\mu_{5}$ of each experiment. As the table suggests, our original method has a significantly higher $\mu_{5}$ than case (i) and case (ii), regardless of the ML model applied. This finding indicates that both the CIMG encoder and the CL-SCL network are important components in Phase Three.
2.2 System Validation via Wet-lab Experiments
2.2.1 Experimental Setup
To obtain 6-(1-methyl-1H-indazol-4-yl), the target product, we consider the following chemical subspace encompassing four key dimensions: reactant, ligand, base, and solvent. Further, considering the typical reaction conditions of Suzuki reactions and the availability of chemicals, the initial chemical subspace, as shown in Figure 4a and 4b, has 10000 possible reactions within it. [^1] The chemical subspace consists of nine reactants (1a-1c react with 2c-2d and 1d-1e react with 2a-2b, leading to 10 combinations), ten ligands, ten bases, and ten solvents, which leads to in 10000 possible reactions.
Figure 4: Overview and procedure of the studied Suzuki-Miyaura reactions. a) The chemical space of the studied Suzuki-Miyaura reaction. b) Structure of phosphorus ligands shown in a). c) All reactions were conducted in Schlenk tube under Ar atmosphere. d) Identification of the product and yield analysis with LC-MS and HPLC. e) The Calculation of the Relative Response Factor of internal standard method; f) Yield analysis of each reaction with HPLC.
After operation and analysis conducted in Phase One and Phase Two, our agent identifies from existing literature that K2CO3 could be the optimal base, and dioxane/H2O could be the most suitable solvent. This insight effectively reduces the unknown variables within our chemical subspace to two dimensions: reactants and ligands. Consequently, the reaction subspace is simplified to 100 viable reaction configurations. Then, Phase Three, using the TT network as the example ML model embedded after the CL-SCL network, makes final recommendations among the narrowed reaction subspace. In a practical application, Phase Three is used only once with fixed CL-SCL and TT networks. However, to comprehensively test our scheme’s robustness in the face of randomness during the training process, we re-examine Phase Three three times, with each time using a different random seed in the training process.
Additionally, we also perform random selections from the original reaction space with 10,000 reactions, which serves as a comparison benchmark to our RAG-AI agent.
2.2.2 Experimental Details and Discussions
We now give a detailed description of our operation in the experiments. Unless otherwise specified, all reagents and solvents were commercially available and used without further purification. All reactions were performed under a positive pressure of argon atmosphere and preliminary monitored by thin-layer chromatography (TLC) while visualized with ultraviolet (UV) light at 254 nm. The product was purified by column chromatography on silica gel 200–300 mesh and identified by high-performance liquid chromatography/mass spectrometry (HPLC-MS), just as depicted in Figure 4d. Fourth, the reaction yield used 4-chloro-6,7-dimethoxyquinoline as an internal standard and was analyzed by HPLC, which was determined by Thermo Scientific Vanquish Core HPLC with a binary pump, a UV detector, an Acclaim 120-C18 (3*150 mm, 3 $\mu$m), using mobile phase A: H2O /B: CH3CN, the flow rate is 0.5 mL/min.
A typical procedure for the synthesis process and yield analysis is as follows: A 20 mL sealed Schlenk tube under argon atmosphere was charged with 6-bromoquinoline (1b, 129 mg, 0.5 mmol), 1-methyl-4-(4,4,5,5-tetramethyl-1,3,2-dioxaborolan-2-yl)-1H-indazole (2d, 103 mg, 0.5 mmol), Pd(OAc)2 (5.66 mg, 0.025 mmol), potassium carbonate (B1, 173 mg, 1.25 mmol) and Amphos (L10, 13.3 mg, 0.05 mmol) in dioxane/H2O (9:1, 5 mL), as shown in Figure 4c. The reaction mixture was heated at 85oC for 16 hours, cooled down to room temperature, and monitored by TLC. Then reaction solution was diluted to 10 mL by methanol, measuring off 0.5 mL diluted solution by pipette, added in 5 mg internal standard, and re-diluted to 10 mL, which was used for HPLC analysis. As the Relative Response Factor (RRF=0.5343) has been determined in Figure 4e and the peak area ratio of product to internal standard is determined in Figure 4f, it is sufficient to calculate the yield is 91.6%. Other reactions were conducted and analyzed following procedures similar to the above steps.
Experimental results are recorded in Table 2. All three experimental batches recommended by Chemist-X reached a $\mu_{5}$ exceeding 90%. In contrast, random sampling within the four-dimension space only resulted in a $\mu_{5}$ around 52%. This stark comparison demonstrates the proficiency of our agent in identifying high-yield reaction conditions amidst a broad chemical reaction space, consistently outperforming the baseline method regardless of possible randomness in the network training process.
Table 2: Results of wet-lab experiments.
Reaction No. Reactant1 Reactant2 Ligand Base Solvent Yield
Experimental Batch #1 1b 2d Amphos K2CO3 Dioxane: H2O = 9:1 0.916
1d 2b dppp K2CO3 Dioxane: H2O = 9:1 0.307
1c 2d PPh3 K2CO3 Dioxane: H2O = 9:1 0.559
1c 2c Xantphos K2CO3 Dioxane: H2O = 9:1 0.741
1a 2d Amphos K2CO3 Dioxane: H2O = 9:1 0.709
Experimental Batch #2 1a 2d PPh3 K2CO3 Dioxane: H2O = 9:1 0.026
1e 2b dppe K2CO3 Dioxane: H2O = 9:1 0.028
1c 2d cataCXium A K2CO3 Dioxane: H2O = 9:1 0.801
1a 2c SPhos K2CO3 Dioxane: H2O = 9:1 0.943
1a 2c Xantphos K2CO3 Dioxane: H2O = 9:1 0.986
Experimental Batch #3 1b 2d PPh3 K2CO3 Dioxane: H2O = 9:1 0.791
1c 2c dppp K2CO3 Dioxane: H2O = 9:1 0.467
1b 2c Xantphos K2CO3 Dioxane: H2O = 9:1 0.836
1d 2a Xphos K2CO3 Dioxane: H2O = 9:1 0.412
1b 2d cataCXium A K2CO3 Dioxane: H2O = 9:1 0.965
Random Batch 1c 2c Xantphos LiOtBu MeCN: H2O = 9:1 0.515
1b 2d SPhos NaOH DME: H2O = 9:1 0.269
1e 2a PPh3 Cs2CO3 DMSO: H2O = 9:1 0.001
- 1e 2b JohnPhos Na2CO3 MeCN: H2O = 9:1 0.351
1a 2c PPh3 K2CO3 Toluene: H2O = 9:1 0.388
3 Method
3.1 Phase One: information retrieval using molecule database memory
Databases like PubChem typically offer API interfaces for user queries, along with comprehensive tutorials that demonstrate the functions of these APIs through examples. In Phase One, we harness the API and its documentation with our innovative TMS-ICL strategy.
Let us start with the ICL technology. Distinct from conventional supervised learning methods, ICL does not require parameter optimization within the network. It allows an LLM to form a “short-term memory” based on specific examples. Given prompts enriched with detailed examples, an LLM can more accurately perform tasks drawing upon the contextual information assimilated from the prompt, which facilitates its stunning performance across various NLP benchmarks.
The efficacy of ICL hinges on the selection of appropriate examples. A highly relevant example can significantly enhance the LLM performance compared to a less related one. Traditional ICL research, such as [33], typically involves manual example selection tailored to a particular task. The ICL process in Phase One seeks to automate this example selection so that the system can adapt to different databases without human intervention. Otherwise, a chemist would need to manually review documentation and choose examples, which would undermine the automation objective of Chemist-X.
We put forth a quantitative method for selecting the document part most relevant to our problem defined in the prompt. Our method begins by dividing the tutorial document into slices with similar lengths. Then, we tokenize the content of a document slice and convert the contextual information within the slice into vector representations using OpenAI’s ADA-002 model [25]. Let us denote the representing vector (also known as embedding) of slice $i$ by $A_{i}$. Concurrently, we process the questioning prompt with the same model and denote its embedding by $B$.
Following this transformation, we compute the cosine similarity between vector $B$ and $A_{i}$ with the following equation
$d_{i}=1.0-\frac{{\sum_{k=1}^{K}}A_{i}(k)\cdot B(k)}{\sqrt{\sum_{k=1}^{K}A_{i}^{2}(k)}\cdot\sqrt{\sum_{k=1}^{K}B^{2}(k)}}$ (1)
where $K$ is the dimension of the embedding space.
An alternative way to measure the similarity is using the L2 similarity written as
$d^{{}^{\prime}}_{i}=\frac{1}{1+\sqrt{{\sum_{k=1}^{K}}\left(A_{i}(k)-B(k)\right)^{2}}}$ (2)
We refer to the document slice with the highest similarity to our question as TMS. It contains the most semantically relevant contexts, and is, therefore, the most suitable slice for our ICL’s example input. In Fig. 1 of the extended data, we provide an example Python code snippet that was automatically generated by GPT during Phase One for illustrative purposes.
3.2 Phase Two: information analysis using online literature memory
Phase Two encompasses two subtasks. The first involves creating web crawlers that extract HTML content from scientific literature websites. We employ web crawlers for data retrieval in this phase, as literature platforms like ChemSpider offer their resources through web interfaces, unlike the APIs utilized in Phase One. The second subtask is to analyze the obtained HTML data and generate a list of potential reaction conditions identified according to the studied literature, which serves as Phase Two’s deliverable.
The development and operation of the web crawlers are guided by ICL prompts. To ensure precise and efficient HTML data extraction, we have devised a specialized prompting scheme for our crawler-based information retrieval. The protocol unfolds as follows. The initial step instructs the agent on setting up the crawling environment, including the installation of necessary Python packages, opening the browser, and logging in to the literature platform with the given username and password. The second step outlines the structure of the target webpage, prompting the agent to emulate human interaction (i.e., clicking and typing) using the crawler. Once the interactions are complete, the third step guides the agent to download the HTML content featuring the search results for the molecule of interest. The final prompt asks the crawler to iterate over the third and fourth steps until all similar molecules identified in Phase One have been searched within literature platforms, with all relevant HTML sources downloaded. Fig. 2 of the extended data shows an example Python realization for an interested molecule.
After downloading these HTML sources, our agent is tasked with composing an HTML analysis function in Python. The example we design for the ICL process should consider complex HTML structures (e.g., multiple reactions are found for the target molecule, and each reaction may consist of multiple possible conditions reported in previous works) to ensure the generated code’s effectiveness under intricate circumstances. This approach contrasts with general HTML analysis methods [17, 18], in which the raw HTML source is directly given to the LLM. Our method is explicitly tailored for the webpages of literature platforms and allows for code reuse in subsequent HTML file analysis. Experimental results in Section 2.1.2 have demonstrated that our approach not only improves the performance of HTML analysis but also minimizes computational resources consumed. We refer readers to Fig. 3 of the extended data for an example code realization completed by GPT-4.
3.3 Phase Three: final recommendation with pre-packaged fingerprint tool
It is important to note that a reaction condition emerging from Phase Two may be incomplete. Although a literature platform could offer comprehensive condition information for numerous chemical reactions, it might inadvertently omit critical details in a reaction, such as reagents, catalysts, solvents, and yield. In other words, detailed reaction data from the original paper might be uncaptured in the literature platform. Chemist-X aims to accelerate chemical reaction designs and facilitate future robot-operated chemical reaction platforms that free human labor by automatically providing complete recommendations with no information loss. In light of this, at the beginning of Phase Three, we fill in the overlooked information with a comprehensive set of potential solutions in our chemical subspace.
After the information-filling process, the agent needs to identify the best possible reaction conditions with the highest expected yield. Previous works have proved that advanced ML algorithms are competent for the RCR task [20, 19, 21, 34]. However, one major problem we noticed in these works is that their reaction encodings, also known as reaction fingerprints, are just straightforward concatenations of encoded molecules. Although these reaction encoding schemes are easy to implement, we want to emphasize the importance of having a reaction fingerprint with additional yield information integrated into it when training an ML model for yield-predicting tasks or RCR tasks targeting optimal yields.
We now illustrate how our model training process is conducted from two aspects: the pre-processing of the reaction data, and the CL-SCL network.
3.3.1 Reaction data processing
we encode all chemicals that appeared in the reaction (e.g., reactants, ligands, base, solvent, etc.) with the CIMG descriptor introduced in [22]. The CIMG descriptor incorporates a comprehensive set of key information about a chemical molecular, specifically nuclear magnetic resonance (NMR) chemical shifts [35] and bond dissociation energy (BDE) [36]. In CIMG, the NMR chemical shifts and BDEs of a molecule are features as vertex and edge of a GNN model. NMR chemical shifts can be naturally represented as properties assigned to the vertices of the graph because they pertain to specific atoms, while BDEs are associated with individual bonds and can thus be captured as edge properties. As a result, GNN can seamlessly integrate chemical information into the molecular descriptors.
With the incorporation of these chemistry features, the CIMG descriptor becomes an informative molecular descriptor and serves as a robust foundation for the reaction embedding, in which we concatenate CIMG descriptors of each chemical component to form the numerical encoding of each reaction. In a Suzuki reaction, for example, we consider the following components: electrophile, nucleophile, catalyst, ligand, base, and solvent, and the concatenated expression of the reaction is written as
${{{x}}}_{i}={{{x}}}_{{\rm{Electrophile}}}\oplus{{{x}}}_{{\rm{Nucleophile}}}\oplus{{{x}}}_{{\rm{Pd\text{ }catalyst}}}\oplus{{{x}}}_{{\rm{Ligand}}}\oplus{{{x}}}_{{\rm{Base}}}\oplus{{{x}}}_{{\rm{Solvent}}}$ (3)
where $\oplus$ denotes concatenation, and elements on the right side of the equation denote the CIMG descriptors of the above-mentioned vital components of a Suzuki reaction.
3.3.2 the CL-SCL network
The training data of the SCL network contains random noise, as the yield of a chemical reaction may be affected by factors including but not limited to measurement errors, material purity, and environmental factors (e.g., humidity and room temperature). These random data noises could depress the effectiveness and robustness of the yield production models. For example, if reaction A has a yield of 90% while reaction B has a yield of 91%, then their yield gap is even smaller than the random noise of the two chemical reactions. In such cases, it is beneficial to treat the yields of these two reactions at the same level so that the ML model can encapsulate the fundamental chemical insights and underlying reaction principles shared among similar reactions. Otherwise, focusing the ML model to learn the regression knowledge and discriminate 91% over 90% only lures the model to learn from random noises. To overcome such problems, this paper categorizes reaction yields, which range from zero to one, into three types: high-yield, medium-yield, and low-yield. We refer to these labels as coarse reaction yield labels, or coarse labels in short.
We now look at how we adapt the SCL network to encode our coarse-label information into the fingerprint. SCL is a learning framework that extends the ideas of self-supervised contrastive learning to the supervised learning setting where label information is available. It was introduced to improve the quality of representations learned by deep learning models, with a particular focus on improving the robustness and generalization of the learned features. The loss function of our SCL network is
$\mathcal{L}_{\text{SCL}}=\sum_{i=1}^{N}\frac{-1}{N_{y_{i}}}\sum_{\begin{subarray}{c}j=1\\ j\neq i\\ \xi(y_{j})=\xi(y_{i})\end{subarray}}^{N}\log\frac{\exp(\textbf{z}_{i}\cdot\textbf{z}_{j}/\tau)}{\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N}\exp(\textbf{z}_{i}\cdot\textbf{z}_{k}/\tau)}$ (4)
where $\mathcal{L}_{SCL}$ denotes the supervised contrastive loss; $N$ is the number of reaction samples; $y_{i}$ is the reaction yield of the $i^{th}$ sample; $\xi(y_{i})$ is the coarse label of $i^{th}$ sample; $N_{y_{i}}$ is the number of reaction samples that have the same coarse label as the sample $i$; $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ are the embeddings of the $i^{th}$ and $j^{th}$samples, respectively; $\tau$ is a temperature scaling parameter that helps to control the separation of the classes.
An intuitive explanation of using the SCL network for our task is as follows. The loss function encourages the model to learn to minimize the distance between embeddings of examples with the same label and maximize the distance between embeddings of examples with different labels. In this way, SCL networks help to learn an embedding space where the representations of reactions are organized based on their coarse labels. This embedding space groups together reactions that have the same coarse label, while also ensuring that reactions that have different coarse labels are well-separated. This separation in the feature space means that the model learns to 1) bring closer the embeddings of samples with the same coarse label and 2) push apart embeddings of samples with different coarse labels. Such a feature space is very helpful in distinguishing the unique features of high-yield, medium-yield, and low-yield reactions.
One important thing we found is that the embedding constructed with the SCL scheme is a very efficient input for a variety of ML models. For example, in image processing research, [37] builds a convolutional neural network after extracting the embedding vector from the SCL network. In this regard, the post-SCL embedding vector is, in essence, a special reaction encoding that can be further processed with different ML models. Therefore, we follow the setting in [29, 30] and define the post-SCL embedding as a reaction fingerprint.
In this paper, we tested three different ML models: random forest [31], tabular transformer [9], and Xgboost [32]. Implementations of the three ML models are packaged within Phase Three’s API function. When calling the API, Chemist-X can select one of the three models for the execution of Phase Three via a user-friendly API interface.
4 Conclusion
This paper introduces Chemist-X, an innovative RAG-AI agent that harnesses the latest research breakthroughs in RAG technology and LLMs to transform the landscape of reaction condition recommendations in chemical synthesis. In general, this paper makes three significant contributions. First, with the “search-analyze-recommend” pattern of expert chemists and the novel reaction fingerprint that integrates yield data, Chemist-X demonstrates remarkable proficiency in automating the RCR task. Second, the use of updated online databases as knowledge sources allows our agent to transcend the limitations of traditional AI systems, which are confined to the knowledge within their training data. Furthermore, the agent’s ability to leverage cutting-edge CAD tools also enables the continuous evolution of the system. Third, the successful application of our RAG-AI agent to the Suzuki reaction example, validated through wet-lab experiments, underscores the practical impact of this research. With the above technical and experimental contributions, this research brings the vision of fully automated chemical discovery one step closer, marking a significant leap in the field of computer-aided chemistry.
5 Extended Data
Extended Data Fig. 1 presents the Code for PubChem API request, which is generated by Chemist-X. The generated code receives SMILES as input and leverages PubChem API to acquire similar molecules. Extended Data Fig. 2 presents the code for the HTML crawler, which is generated by Chemist-X. The generated code can simulate the following steps: clicking the relevant button, inputting the SMILES of molecules into the search bar, downloading the final HTML, and returning to the website’s homepage. Extended Data Fig. 3 presents the code for the HTML analyzer, which is generated by Chemist-X. The generated code includes the following steps: HTML parsing, the extraction of useful information, and the saving of useful information.